OpenStudy (anonymous):

PLEASE HELP Your friend hands you a graph of the performance statistics of the latest powerboat to be produced. He says, “I know this graph is f(x) = –2(x – 4)3 – 3 but I can’t remember how it is related to the graph of x3.” Explain to your friend how the graph f(x) is a translation of the graph x3.

OpenStudy (anonymous):

@zepdrix @Zarkon @myininaya @e.mccormick @thomaster @robtobey @jigglypuff314

jigglypuff314 (jigglypuff314):

y = a(x-h)^3 + k were h is how much it was moved left to right k is how much up or down and a is how much stretched/squashed

OpenStudy (anonymous):

I don't understand how that relates to the question...@jigglypuff314

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

the original parent graph is y = x^3 = 1(x-0)^3+0 as in it was not moved or stretched in comparison to -2(x-4)^3-3

OpenStudy (anonymous):

I'm sorry..I don't understand.

jigglypuff314 (jigglypuff314):

if y = x^3 is the original graph, then -2(x-4)^3-3 is just that x^3 has been flipped upside down (because of the negative in front) the 2 means it was stretched by a factor of 2 the 4 means that it was shifted 4 to the right the -3 means that it was shifted 3 down

OpenStudy (anonymous):

So how could I answer the question in complete sentences?

jigglypuff314 (jigglypuff314):

well... I'm not good with grammer and complete sentence stuff. so I don't think I can help you with that, sorry :/ I gave all the information you would need to answer the question and another way to explain it is by graphing both graphs and comparing that way